# Mamba-FCS: Joint Spatio-Frequency Feature Fusion, Change-Guided Attention, and SeK Loss for Enhanced Semantic Change Detection in Remote Sensing

### v1

@@ V-B Effect of FFT2 Branch
Removing the FFT2 Branch results in performance drops of 1.47%, 0.91%, 1.26%, and 0.76% in SeK, mIoU, $F_{scd}$, and OA, respectively, as shown in Table I. This underscores the FFT2 Branch’s critical role in enhancing model performance. In Figure 6, rows (a) and (b) demonstrate the FFT2 Branch’s ability to recover sub-pixel edges and suppress hallucinatory changes, which are prevalent without the FFT2 Branch. In rows (c) and (d), under challenging conditions such as heavy shadows and low contrast, the FFT2 Branch effectively detects transitions from low vegetation to ground. Rows (e) and (f) show FFT2 branch reduces boundary erosion and false changes in unchanged structures, alongside improved detection of changes from buildings to non-vegetated surfaces and reduced false detections of buildings on ground in row (f).
@@
V-C Effect of Change-Guided Attention (CGA)
Table I shows that removing the CGA module leads to performance decreases of 1.43%, 0.14%, 2.17%, and 0.54% in SeK, mIoU, $F_{scd}$, and OA, respectively. Figure 7 further illustrates the pivotal role of CGA, which interconnects the three decoders in our architecture ($\mathcal{F}_{BCD}$, $\mathcal{F}^{T_{1}}_{SCD}$, and $\mathcal{F}^{T_{2}}_{SCD}$). Rows (a) to (d) demonstrate CGA’s ability to improve change map fidelity, and suppress false changes around changed areas. Rows (e) and (f) highlight CGA’s benefits for both BCD and SCD tasks, showing that without CGA, independent operation of decoders leads to fragmented change masks and mislabeled semantics at boundaries. By facilitating information exchange among decoders, CGA enhances coordinated learning, directly contributing to the quantitative gains observed in Table I.
@@ V-C Effect of Change-Guided Attention (CGA)
Figure 8: The qualitative impact of omitting the Separated Kappa (SeK). Columns (from left to right) display bi-temporal inputs, ground truth, results without SeK, and the full model. Red boxes highlight regions where omitting SeK increases false semantic labels, while including SeK reduces artifacts.
V-D Effect of SeK Loss
@@
Removing the SeK Loss during training results in performance drops of 0.79%, 0.40%, 0.69%, and 0.19% in SeK, mIoU, $F_{scd}$, and OA, respectively, as shown in Table I. The most significant drop occurs in the SeK metric, which emphasizes class-balanced semantic change identification, confirming that SeK Loss primarily enhances semantic class accuracy. Figure 8 provides qualitative evidence supporting this observation. Rows (a) and (c) illustrate how SeK Loss reduces false occurrences of low vegetation on ground. Similarly, rows (d) and (f) show its effectiveness in mitigating false occurrences of trees on low vegetation. Row (e) demonstrates a complete elimination of false occurrences of water. Notably, SeK Loss enhances differentiation between changed and unchanged regions across all rows, slightly improving change map accuracy.
@@
In summary, SeK Loss strengthens semantic class identification by prioritizing class-balanced change detection, reducing false changes and improving accuracy, as evidenced by quantitative metrics in Table I and qualitative results in Figure 8.
V-E Comparative Experiments
@@
To assess our model’s performance, we benchmark against state‐of‐the‐art methods for SCD. The CNN‐based competitors include HRSCD variants (S2–S4) [39], ChangeMask [38], SSCD-1 and Bi-SRNet [29], and TED [42]. Transformer‐based baselines are SMNet [62] and ScanNet [42]. We also compare our model with Mamba-based baselines such as ChangeMamba [28].
V-E1 Quantitative and Qualitative Results
@@
We present quantitative results for the SECOND dataset and for the Landsat-SCD dataset in Table II. For clarity, SOTA competitors are categorized into three families: (i) CNN-based models, (ii) Transformer-based models, and (iii) Mamba-based models. Qualitative comparisons are provided through segmentation maps in Fig. 9 for the SECOND dataset and Fig. 10 for the Landsat-SCD dataset.
@@
Figure 9: Qualitative comparison on the test split of SECOND dataset. Columns show (i) bi-temporal inputs, (ii) ground truth, and predictions from HRSCD-S4, Bi-SRNet, ScanNet, ChangeMamba, and our Mamba-FCS.
@@
These results are corroborated by the qualitative analysis in Fig. 9. In rows (a), (b), (d), (h), and (i), HRSCD-str4 and Bi-SRNet consistently fail to identify the correct semantic class. While ScanNet and ChangeMamba accurately detect semantic labels in these cases, Mamba-FCS outperforms them by significantly reducing false change detections and enhancing segmentation quality, closely aligning with the GT.
TABLE II: Semantic change-detection results on the SECOND and Landsat-SCD datasets (higher is better). For each dataset, best, second-best and third-best scores are highlighted in red, blue and green, respectively. Standard train-test splits were used for both datasets.
@@
Method OA F1 mIoU SeK OA $F_{\text{scd}}$ mIoU SeK
@@
Figure 10: Qualitative comparison on the test split of Landsat-SCD dataset. Columns display (i) bi-temporal inputs, (ii) ground truth, and predictions from HRSCD-S4, Bi-SRNet, SCanNet, ChangeMamba, and Mamba-FCS. Red boxes highlight regions of improved performance for reader convenience.
These findings are supported by the qualitative analysis in Fig. 10. In rows (a) and (b), OurModel captures water regions more accurately than competing models. Similarly, in rows (c) to (j), OurModel demonstrates superior class detection, minimizes false change alarms, and produces segmentation results that closely match the GT. For readers convenience we highlight the areas of improvements in Red.
Figure 11: Dataset-level from-to confusion matrices for test split of SECOND (first row) and Landsat-SCD (second row). Cell colors indicate the percentage of changed pixels. Columns (from left to right) display ground truth, SCanNet, ChangeMamba, and Mamba-FCS. Mamba-FCS accurately captures ground-truth distributions for dominant transitions while minimizing noise in rare class predictions.
V-E2 Change Analysis
To evaluate the efficacy of each model in accurately capturing both dominant and rare class transitions, we constructed dataset-level ”from $\rightarrow$ to” confusion matrices by aggregating all changed pixels across the official test splits of the SECOND and LandSat-SCD benchmarks, as presented in Figure 11.
In the SECOND dataset, GT identifies major transitions as ground $\rightarrow$ building (32.01%), low-vegetation $\rightarrow$ ground (14.60%), and low-vegetation $\rightarrow$ building (11.59%). Our proposed Mamba-FCS closely matches these key transitions, achieving corresponding values of 34.50%, 14.06%, and 14.65%, respectively. The deviations from GT are consistently within 2.5%, except for low-vegetation $\rightarrow$ building, which differs by 3.06%. Conversely, SCanNet significantly overestimates the ground $\rightarrow$ building transition, predicting it at 38.65%, while ChangeMamba marginally shifts the low-vegetation $\rightarrow$ building transition to 13.14%. For rare transitions involving classes such as tree, water, and playground (each less than 3% occurrence in GT), Mamba-FCS maintains noise levels below 0.9%. Notably, Mamba-FCS predicts water $\rightarrow$ building at 0.29% and playground $\rightarrow$ ground at 0.26%, significantly outperforming SCanNet (8.6% overall noise) and ChangeMamba (8.1% overall noise), with Mamba-FCS maintaining a total noise of only 4.2% across the ten least frequent transitions.
In the LandSat-SCD dataset, GT predominantly includes desert $\rightarrow$ farmland (62.40%), desert $\rightarrow$ water (11.95%), and farmland $\rightarrow$ desert (9.13%) transitions. Our Mamba-FCS accurately aligns with these major transitions, reporting values of 62.98%, 11.96%, and 8.34%, respectively, with minimal deviations (all within 0.79%). In contrast, SCanNet overpredicts the desert $\rightarrow$ farmland transition at 68.10%, and ChangeMamba slightly overrepresents the desert $\rightarrow$ building transition (4.40% versus GT’s 3.84%). For infrequent transitions such as building $\rightarrow$ water and water $\rightarrow$ building (each less than 1.5% occurrence in GT), Mamba-FCS effectively controls noise levels at 0.20%, outperforming ChangeMamba (0.44%) and SCanNet (0.50%). Across all rare transitions, the cumulative noise introduced by Mamba-FCS is 2.1%, notably lower than the 3.7% from ChangeMamba and 3.9% from SCanNet.
These results demonstrate that the proposed Mamba-FCS robustly captures major land cover transitions while effectively mitigating noise from minor, infrequent class changes, underscoring its suitability and reliability for environmental monitoring and urban growth analysis tasks.
@@ VI Conclusions
In this paper, we presented Mamba-FCS, a SCD framework specifically designed to effectively address the challenges of capturing long-range contextual dependencies and detecting subtle semantic transitions in high-resolution remote sensing imagery. Our approach combines the efficiency and global contextual modeling capabilities of VMamba, a linear-complexity state-space backbone, with an novel Joint Spatio-Frequency Feature Fusion strategy that incorporates log-amplitude frequency domain features to mitigate illumination-related artifacts and enhance fine-grained boundary detection. Additionally, we integrated a lightweight Change-Guided Attention (CGA) mechanism, which aligns semantic prediction heads with binary change cues, thereby improving semantic accuracy. Furthermore, we proposed the Separated Kappa(SeK) loss function, repurposing the established SCD evaluation metric into an effective training objective, particularly benefiting minority class transitions.
Extensive experiments on the SECOND and Landsat-SCD benchmarks demonstrate that Mamba-FCS outperforms recent methods across multiple metrics, including overall accuracy, F-score, mean Intersection over Union (mIoU), and SeK. Notably, our framework excels in detecting rare and subtle semantic transitions, achieving reduced false positives and improved class-wise accuracy. These results underscore the efficacy of Mamba-FCS in addressing complex SCD challenges in remote sensing.

### v2

@@ V-A Ablation Studies on Key Contributions
Quantitative results, are reported in Table II. Qualitative analyses on SECOND dataset for the FFT2 branch, CGA, and Sek Inspired Loss are presented in Figures 7, 8, and 9, respectively. Each figure is structured with four columns. The first displays input bi-temporal images, the second shows the Ground Truth (GT), the third presents the model’s output without the ablated component, and the fourth depicts the full model’s output. Red boxes highlight regions where each component significantly enhances performance. The results in Table II and the qualitative visualizations collectively demonstrate that each component plays a valuable role in improving the model’s performance for SCD.
@@ V-B Effect of FFT2 Branch
Removing the FFT2 Branch results in performance drops for both benchmarks, as shown in Table II. This underscores the FFT2 Branch’s critical role in enhancing model performance. In Figure 7, rows (a1) and (a2) demonstrate the FFT2 Branch’s ability to recover sub-pixel edges and suppress hallucinatory changes, which are prevalent without the FFT2 Branch. In rows (b1) and (b2), under challenging conditions such as heavy shadows and low contrast, the FFT2 Branch effectively detects transitions from low vegetation to ground. Rows (c1) and (c2) show FFT2 branch reduces boundary erosion and false changes in unchanged structures, alongside improved detection of changes from buildings to non-vegetated surfaces and reduced false detections of buildings on ground in row (f).
@@
V-C Effect of Change-Guided Attention (CGA)
Table II shows that removing the CGA module leads to performance decreases of 1.43%, 0.14%, 2.17%, and 0.54% in SeK, mIoU, $F_{\text{scd}}$, and OA, respectively, for the SECOND dataset, and 3.78%, 0.92%, 2.01%, and 0.50% for the Landsat-SCD dataset. Figure 8 further illustrates the pivotal role of CGA, which interconnects the three decoders in our architecture ($\mathcal{F}_{BCD}$, $\mathcal{F}^{T_{1}}_{SCD}$, and $\mathcal{F}^{T_{2}}_{SCD}$). Rows (a1) to (b2) demonstrate CGA’s ability to improve change map fidelity, and suppress false changes around changed areas. Rows (c1) and (c2) highlight CGA’s benefits for both BCD and SCD tasks, showing that without CGA, independent operation of decoders leads to fragmented change masks and mislabeled semantics at boundaries. By facilitating information exchange among decoders, CGA enhances coordinated learning, directly contributing to the quantitative gains observed in Table II.
@@ V-C Effect of Change-Guided Attention (CGA)
Figure 9: The qualitative impact of omitting the Separated Kappa (SeK) Loss. Columns (from left to right) display bi-temporal inputs, ground truth, results without SeK, and the full model. Red boxes highlight regions where omitting SeK Loss increases false semantic labels, while including SeK reduces artifacts.
V-D Effect of SeK Inspired Loss
@@
Removing the SeK-inspired loss during training results in performance drops of 0.79%, 0.40%, 0.69%, and 0.19% in SeK, mIoU, $F_{\text{scd}}$, and OA, respectively, for the SECOND dataset, and 3.70%, 1.29%, 2.51%, and 0.48% for the Landsat-SCD dataset, as shown in Table II. The most significant drop occurs in the SeK metric, which emphasizes class-balanced semantic change identification, confirming that Sek Inspired Loss primarily enhances semantic class accuracy. Figure 9 provides qualitative evidence supporting this observation. Rows (a1) and (b1) illustrate how Sek Inspired Loss reduces false occurrences of low vegetation on ground. Similarly, rows (b2) and (c2) show its effectiveness in mitigating false occurrences of trees on low vegetation. Row (c1) demonstrates a complete elimination of false occurrences of water. Notably, Sek Inspired Loss enhances differentiation between changed and unchanged regions across all rows, slightly improving change map accuracy.
@@
In summary, Sek Inspired Loss strengthens semantic class identification by prioritizing class-balanced change detection, reducing false changes and improving accuracy, as evidenced by quantitative metrics in Table II and qualitative results in Figure 9.
V-E Combined Effect of CGA and SeK Inspired Loss
Table II shows that removing both the CGA module and Sek Inspired Loss results in the largest performance degradation, confirming their synergistic effect.
The Sek Inspired Loss directly supervises the semantic decoders, rewarding accurate semantics within change regions, but does not directly influence the Binary Change Decoder. The CGA module, however, connects the Binary Change Decoder to the Semantic Map Decoders, enabling indirect propagation of Sek Inspired Loss gradients to the Binary Change Decoder. Consequently, with CGA enabled, the Binary Change Decoder benefits from Sek Inspired Loss supervision, enhancing both change detection and semantic consistency. This interplay explains the significant performance drop when both components are removed, validating their combined use in Mamba-FCS.
V-F Loss Function Analysis
To justify our loss design for SCD under severe class imbalance, we conduct two complementary ablations on the SECOND dataset.
First, we compare three additional imbalance handling terms Focal[27], Dice[40] and the proposed SeK inspired term while keeping the baseline supervision and the Mamba-FCS architecture fixed. As summarised in Table III, the SeK inspired term achieves higher SeK and mIoU than both Focal and Dice losses, while OA and $F_{\text{scd}}$ remain very similar, indicating a more effective focus on under-represented semantic transitions.
TABLE III: Effect of different additional loss terms for handling class imbalance on the SECOND dataset.
Additional term Loss components SECOND
Focal Dice SeK OA (%) $F_{\text{scd}}$ (%) mIoU (%) SeK (%)
Focal $\checkmark$ 88.23 65.36 73.90 24.90
Dice $\checkmark$ 88.42 64.93 73.86 24.81
SeK-inspired (ours) $\checkmark$ 88.62 65.78 74.07 25.50
Second, we study the weighting coefficients $\lambda_{1}$ and $\lambda_{2}$ in the SeK-inspired loss. Table IV shows that extreme or unbalanced weights reduce SeK and mIoU, whereas the configuration $\lambda_{1}=0.15$ and $\lambda_{2}=0.30$ provided a reasonable balance across metrics and is used in our experiments to achieve the best performance.
TABLE IV: Sensitivity of the SeK-inspired loss to different choices of $\lambda_{1}$ and $\lambda_{2}$ on the SECOND dataset
$\lambda_{1}$ $\lambda_{2}$ OA (%) $F_{\text{scd}}$ (%) mIoU (%) SeK (%)
1.00 1.00 88.36 64.94 73.85 24.84
0.50 1.00 88.58 65.15 73.98 25.03
0.30 0.15 88.06 64.77 73.32 24.25
0.15 0.30 88.62 65.78 74.07 25.50
V-G Comparative Experiments
@@
To assess our model’s performance, we benchmark against SOTA methods for SCD. The CNN‐based competitors include HRSCD variants (S2–S4) [16], ChangeMask [58], SSCD-1 and Bi-SRNet [17], and TED [18]. Transformer‐based baselines are SMNet [36] and ScanNet [18]. We also compare our model with Mamba-based baselines such as ChangeMamba [12].
V-G1 Quantitative and Qualitative Results
@@
We present quantitative results for the SECOND dataset and for the Landsat-SCD dataset in Table V. For clarity, SOTA competitors are categorized into three families: (i) CNN-based models, (ii) Transformer-based models, and (iii) Mamba-based models. Qualitative comparisons are provided through segmentation maps in Fig. 10 for the SECOND dataset and Fig. 11 for the Landsat-SCD dataset.
@@
Figure 10: Qualitative comparison on the test split of SECOND dataset. Columns show (i) bi-temporal inputs, (ii) ground truth, and predictions from (iii) HRSCD-S4, (iv) Bi-SRNet, (v) SCanNet, (vi) ChangeMamba, and (vii) Mamba-FCS.
@@
These results are corroborated by the qualitative analysis in Fig. 10. In rows (a1), (a2), (b2), (d1), and (e1), HRSCD-str4 and Bi-SRNet consistently fail to identify the correct semantic class.nWhile ScanNet and ChangeMamba generally recover the correct semantic classes, their predictions still exhibit scattered false change blobs around building footprints and vegetation, particularly along boundaries. In contrast, Mamba-FCS produces more compact building and playground regions, suppresses false changes in surrounding low-vegetation and tree areas, and better preserves unchanged structures. This systematic reduction of boundary noise and rare false transitions is consistent with the observed improvements in $F_{scd}$ and SeK, and is aligned with the design of the FFT2 branch and CGA, which respectively sharpen high-frequency edges and enforce coherence between the binary change and semantic decoders.
TABLE V: Semantic change-detection results on the SECOND and Landsat-SCD datasets (higher is better). For each dataset, best, second-best and third-best scores are highlighted in red, blue and green, respectively. Standard train-test splits were used for both datasets.
@@
Method OA $F_{\text{scd}}$ mIoU SeK OA $F_{\text{scd}}$ mIoU SeK
@@
Figure 11: Qualitative comparison on the test split of Landsat-SCD dataset. Columns display (i) bi-temporal inputs, (ii) ground truth, and predictions from (iii) HRSCD-S4, (iv) Bi-SRNet, (v) SCanNet, (vi) ChangeMamba, and (vii) Mamba-FCS. Red boxes highlight regions where performance improvement is clearly visible.
These findings are supported by the qualitative analysis in Fig. 11. In rows (a1) and (a2), Mamba-FCS captures water regions more accurately than competing models. Similarly, in rows (b1) to (e2), Mamba-FCS substantially reduces these false alarms, yielding cleaner desert extents and more coherent farmland and water patches, particularly along transition fronts. This behaviour matches the quantitative gains in SeK, where Mamba-FCS better preserves the distribution of major transitions while reducing noise for rare ones. For readers’ convenience, we highlight regions with noticeable improvements in red.
Figure 12: Dataset-level from–to confusion matrices and corresponding error maps for the SECOND test split. The top row shows the percentage of changed pixels for each land-cover transition for gronud truth (a), SCanNet (b1), ChangeMamba (c1), and Mamba-FCS (d1); the bottom row (b2–d2) shows signed differences.
V-G2 Change Analysis
To provide a more interpretable view of how each model behaves on different land-cover transitions, we plot dataset-level “from $\rightarrow$ to” confusion matrices and corresponding signed error maps for the SECOND and LandSat-SCD test splits in Figures 12 and 13. Each confusion matrix in the top row summarises the percentage of changed pixels moving from a source class (rows) to a target class (columns) for the ground truth and for the predictions of SCanNet, ChangeMamba, and Mamba-FCS. The error maps in the bottom row then visualise, for every transition, the difference between the model prediction and ground truth, where red cells indicate overestimation and blue cells indicate underestimation.
On the SECOND dataset (Figure 12), the ground-truth matrix is dominated by transitions such as ground $\rightarrow$ building and low vegetation $\rightarrow$ ground/building, while transitions involving tree, water, and playground are comparatively rare. SCanNet’s error map (b2) shows a strong positive bias on the ground $\rightarrow$ building entry and noticeable negative deviations on several other entries, revealing a tendency to overestimate urban expansion at the expense of other change patterns. ChangeMamba (c2) reduces some of these extremes but still exhibits distinct red/blue patches, particularly on transitions towards the building class. In contrast, the error map of Mamba-FCS (d2) is largely composed of light colours with only small residual deviations, indicating that the predicted distribution of land-cover conversions closely follows the ground truth. Rare transitions in the lower-right region of the matrix remain close to zero in both value and error, showing that Mamba-FCS avoids introducing spurious changes for infrequent categories.
Figure 13: Dataset-level from–to confusion matrices and corresponding error maps for the Landsat-SCD test split. The top row shows the percentage of changed pixels for each land-cover transition for ground truth (a), SCanNet (b1), ChangeMamba (c1), and Mamba-FCS (d1); the bottom row (b2–d2) shows signed differences
A similar trend appears on the LandSat-SCD dataset (Figure 13). Here, the ground truth is dominated by transitions such as desert $\rightarrow$ farmland and desert $\rightarrow$ water, with other conversions occurring much less frequently. SCanNet’s error map (b2) contains pronounced red regions on desert $\rightarrow$ farmland and related entries, indicating systematic overestimation of agricultural expansion, while ChangeMamba (c2) still shows visible positive and negative patches across several desert-related transitions. In contrast, the error map of Mamba-FCS (d2) is almost uniformly close to zero, with only mild residual deviations on the dominant transitions and very small errors on rare ones, such as conversions between building and water.
Overall, these confusion matrices and signed error maps confirm that Mamba-FCS not only improves scalar metrics such as SeK and mIoU, but also reduces systematic biases in specific land-cover conversions and better preserves the empirical distribution of both common and rare transitions. This behaviour is especially desirable for environmental monitoring and urban-growth analysis, where over- or under-estimating particular transitions (e.g., urbanisation or farmland expansion) could lead to misleading conclusions about the underlying dynamics.
V-G3 Computational Cost and Efficiency
The table VI compares the computational cost for Mamba Based methods with transformer based models. Although Mamba-FCS has the largest number of parameters among the compared methods, its computational cost remains on par with the Transformer baseline ScanNet (263.15 vs. 264.95 GFLOPs) despite having roughly 7 times more parameters, and only moderately higher than ChangeMamba (211.55 GFLOPs) while using more than 2 times the parameters. This decoupling between parameter count and FLOP cost stems from the linear-time state-space operations in VMamba, which avoid the quadratic scaling of self-attention with spatial resolution.
In practice, this means that Mamba-FCS can afford a higher-capacity backbone and richer decoders without incurring a proportional increase in inference-time computation, making it more suitable for large-scale, high-resolution SCD deployments.
TABLE VI: Comparison of best performing SCD models in computational cost (input: 512 $\times$ 512 two-epoch RGB).
Method Params (M) GFLOPs
Transformer-based models
ScanNet [18] 27.90 264.95
Mamba-based models
ChangeMamba [12] 89.99 211.55
Mamba-FCS 189.54 263.15
@@ VI Conclusions
In this paper, we presented Mamba-FCS, a SCD framework specifically designed to effectively address the challenges of capturing long-range contextual dependencies and detecting subtle semantic transitions in high-resolution remote sensing imagery.
Our approach combines the efficiency and global contextual modeling capabilities of VMamba, a linear-complexity visual state-space backbone, with a novel Joint Spatio-Frequency Feature Fusion strategy that incorporates log-amplitude frequency domain features to mitigate illumination-related artifacts and enhance fine-grained boundary detection. Additionally, we integrated a lightweight Change-Guided Attention (CGA) mechanism, which aligns semantic prediction heads with binary change cues, thereby improving semantic accuracy. Furthermore, we proposed the Separated Kappa(SeK) inspired loss function, re-purposing the established SCD evaluation metric into an effective training objective, particularly benefiting minority class transitions.
Extensive experiments on the SECOND and Landsat-SCD benchmarks demonstrate that our method outperforms recent SOTA approaches across multiple metrics, including overall accuracy, F1-score, mean Intersection over Union (mIoU), and Separability and Kappa (SeK). Specifically, our method achieves a SeK score of 25.50% on the SECOND dataset and 60.26% on the Landsat-SCD dataset. These results highlight its capability in detecting rare and subtle semantic transitions, with reduced false positives and enhanced class-wise accuracy. These results underscore the efficacy of Mamba-FCS in addressing complex SCD challenges in remote sensing.
However, the SeK-inspired loss is defined under fully supervised, pixel-wise annotations, and its behavior under label noise or in weakly/semi-supervised regimes has yet to be explored. Future work will therefore focus on developing lighter or distilled variants of Mamba-FCS for efficient deployment, extending the joint spatio–frequency fusion to multi-modal inputs, and adapting the SeK-guided training strategy to semi-supervised and domain-adaptive SCD scenarios.